OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point. y sin 2x = x cos 2y, (π/4, π/2)

OpenStudy (cwrw238):

dy/dx * sin2x + y* 2 cos2x = cos2y + x*-2sin2y* dy/dx find dy/dx in terms of x and y this will give the slope of the line plug in the given values of x and y and use the general form y-y1 = m(x -x1) to get the required equation

OpenStudy (anonymous):

I'm having trouble with finding dy/dx. I've been trying to understand implicit differentiation but it's not working out

OpenStudy (cwrw238):

you take y to be a function of x eg x^2 y = y^2 use product rule to differentiate x^2 y x^2 * dy/dx + y * 2 = 2y*dy/dx to differentiate y^2 you use the chain rule

OpenStudy (cwrw238):

sorry but i gotta go hope the above helps